# Evaluating Foundation Models’ 3D Understanding Through Multi-View Correspondence Analysis

### v1

@@ E.1 Reproduction
E.2 Experiment A
In experiment A, we assess model generalization across view-angle difficulties using per-class performance curves. In the figures below, the bars represent mean mIoU for unseen validation bins, and dots mark reference viewpoints. Performance typically declines as viewpoint distance increases from the reference bins, consistent with the trend in Figure 2. Across viewpoint difficulties, DINO, DINOv2, and CLIP consistently outperform all other models. C-RADIO follows closely, while SigLIP2 and especially TIPS degrade substantially under extreme viewpoints (with TIPS showing the weakest overall robustness). DINOv3 generally outperforms all models; however, this does not hold for the coat rack and guitar stand classes, where, in the Extreme difficulty scenario, it becomes one of the worst-performing models. This drop may be influenced by the reduced per-bin image counts (as noted in Table 6), making the class results less stable.
While the other models exhibit consistent viewpoint robustness, the VGGT-1B model behaves markedly differently. Its mIoU remains low across all classes and difficulty levels, typically below 0.4, with only minor variation between bins. This is expected given the architectural mismatch with Hummingbird: VGGT is a geometry-grounded, multi-view transformer, and in our setting, it receives only a single view and must rely on adapted aggregator tokens rather than native patch embeddings. Under these conditions, its predictions often collapse toward background-heavy masks, yielding high background mIoU but near-zero scores for small or fine-grained objects. As a result, its degradation curve appears shallow, not because the model is stable, but because performance stays uniformly low in this adapted, single-view configuration.
Across all models, the background class performs best; other classes that perform well are durian, broccoli, and strings. These tend to generalize above average under unseen viewpoints. In contrast, bed, coat rack, guitar stand, and sink perform worst (the largest negative generalization gaps). For the sink class, DINOv3 significantly outperforms all other models. For bed, the drop is explained by poor ground-truth annotations that likely reduce model scores (as explored in Appendix E.2.1).
Figure 5: Experiment A: background. Viewpoint generalization for the background class across four training difficulty settings: (a) easy, (b) medium, (c) hard, and (d) extreme. All models achieve near-perfect consistency across bins, indicating that the background segmentation task remains invariant to viewpoint changes. Since VGGT scores highly only on background while remaining low on all object classes, this discrepancy indicates that the model may over-predict background regions, leading to inflated background mIoU despite poor object segmentation.
@@ E.2 Experiment A
Figure 7: Experiment A: bed. The bed class scores far below the all-class baseline, with it’s mIoU scores remaining low even for strong models. This degradation is largely explained by systematically incorrect ground-truth masks, as seen in Figures 2226 we analyze in Appendix E.2.1. The figure reflects this annotation-driven performance drop.
Figure 8: Experiment A: toy cat. The toy cat performance is very similar to the dataset-wide average observed in Figure 2, with relative differences near zero. Strong models (CLIP, DINO) achieve high mIoU (0.77–0.84) and accuracy drops consistently across difficulty settings.
@@
Figure 10: Experiment A: microwave. This class underperforms relative to Figure 2, with the largest drops appearing under Hard and Extreme difficulty. Model behaviour diverges notably: DINOv3, DINOv2 and CLIP maintain comparatively strong performance, whereas DINO shows a mild decline. TIPS and SigLIP2 appear to struggle when segmenting the microwave category and degrade sharply under the Extreme condition, with a pronounced collapse beyond the 45 bin. Despite these weaker generalization patterns, the overall degradation across viewpoint bins follows the expected trend.
@@
Figure 12: Experiment A: toy dragon. This class exhibits mildly negative gaps, in particular C-RADIO is always lower than TIPS (when comparing to Figure 2). As a whole though, its performance trends mirror the general degradation observed of weaker models struggling as viewpoint distance increases.
Figure 13: Experiment A: coat rack. The coat rack is the worst-performing class overall, with mIoU remaining below 0.7 across all validation bins. VGGT performs particularly poorly, often producing near-zero mIoU. This likely indicates a failure to localize the object rather than viewpoint sensitivity. As seen in Figure 4, the object’s structure is extremely thin, which likely makes it highly sensitive to viewpoint changes and contributes to its weak generalization. It remains an open question whether this difficulty is intrinsic to the geometry or arises from dataset-specific factors. This could be an area of further investigation.
Figure 14: Experiment A: guitar stand. The Guitar stand class scores well below the baseline for all models and difficulties. Degradation is steep: DINOv3 consistently struggles with the 30°validation bin, performs very poorly in the Extreme task, and generally appears to have a large negative gaps (drops by more than 0.2 mIoU) when compared to the “all-class” baseline. VGGT once again produces near-zero mIoU. Like the coat rack, the thin structure likely contributes to poor performance. However, exact cause remains unclear and is a point for future investigation.
@@
Figure 16: Experiment A: toilet. The toilet class consistently performs above the overall trend in Figure 2, with C-RADIO showing the strongest boost, achieving increases of nearly 0.2 mIoU across difficulty levels.
Figure 17: Experiment A: sink. The sink class is one one of the worst three class performances in the dataset. DINOv3 is the only model which performs welll in this category. All other models show large negative gaps relative to Figure 2, with SigLIP2 already underperforming in the easy setting by roughly 0.2 mIoU. Absolute scores generally remain low across all difficulty levels, and the degradation with viewpoint variation is steep. What geometric or appearance factors make sinks perform poorly is not immediately obvious, suggesting that sink segmentation is a broader failure mode worth investigating.
Figure 18: Experiment A: strings. The strings class generalises reliably across all models, with performance declining smoothly as viewpoint difficulty increases yet remaining consistently above the dataset average. Since this category spans multiple instruments (e.g. guitar, ukulele, guzheng), future work could investigate which visual attributes they share which lead to strong viewpoint stability.
@@
Figure 20: Experiment A: durian.. Durian is among the best-performing classes, showing high absolute mIoU (0.78–0.94) across most models and all difficulty levels. Despite VGGT underperforming other models, its performance in this category is among its best, which reinforces this class’s easy difficulty. This robustness is somewhat surprising: although the durian has a rich surface texture, its overall geometry is nearly spherical. It remains unclear whether the stable segmentation under viewpoint shifts stems from its distinctive visual signature or from contextual cues in the surrounding scene, such as the curtains highlighted in Figure 4.
E.2.1 Bed category ground truth analysis
@@
E.3 Experiment B
As shown in Figure 3 and Figure 27, most models exhibit a gradual performance decline as the distance of the validation bin from the training bin (0) increases. Visually, only C-RADIOv2 (ViT-B/16-CPE), TIPS (ViT-B/14-HR) and VGGT exhibit sharp performance drops with increasing validation angles, indicating limited generalization capacity under extreme viewpoint shifts.
Figure 27: Absolute mIoU under viewpoint shifts. Each curve shows the mIoU of a model evaluated on increasing validation bins, measured directly (without normalization to the 0 baseline). The horizontal axis represents viewpoint angle relative to the training bin (0), and the vertical axis shows the segmentation performance via mIoU. All models show a gradual decrease as the angular distance grows, reflecting reduced correspondence between memory and query views. DINOv2 has the most consistent generalization across bins, maintaining a high mIoU across all bins. DINOv3 maintains the highest mIoU, but has an unexpected peak at bin 60. Further research could explore whether this is attributed to DINOv3’s architecture (e.g. its use of register tokens and its more complex pretraining objective), or if it’s a result of the reduced per-bin image counts. By contrast, TIPS displays the steepest decline (dropping by more than 0.2 between 15 and 75), confirming its reduced robustness under large viewpoint changes.
E.4 Experiment C
In experiment C, we evaluate the impact of memory bank size on model performance. Table 10 reports the absolute mIoU gains per difficulty level when increasing the memory bank for enteries (a) from 320k to 640k, (b) from 640k to 1,024k, and (c) from 320k to 1,024k. This highlights how memory sensitivity varies across model architectures. On the other hand, Table 11 outlines the computational trade-offs. We report the runtime and memory usage associated with each model under both memory configurations. This provides context for the computational cost underlying the performance results.
Table 10: Performance gains from memory. We report absolute mIoU improvements when increasing memory from (a) 320k to 640k, (b) 640k to 1,024k, and (c) 320k to 1,024k. Values are computed as differences from Table 4, with the final column showing the average gain across all four difficulty levels.
@@ E.4 Experiment C
CLIP ViT-B/16 0.016 0.015 0.017 0.013 0.01525
DINO ViT-B/16 0.027 0.025 0.024 0.020 0.02400
DINOv2 ViT-B/14 0.016 0.014 0.014 0.012 0.01400
DINOv3 ViT-B/16 0.011 0.009 0.005 0.002 0.00675
C-RADIOv2 ViT-B/16-CPE 0.041 0.036 0.033 0.024 0.03350
SigLIP 2 B/16-512 0.035 0.030 0.028 0.023 0.02900
TIPS ViT-B/14-HR 0.044 0.038 0.032 0.016 0.03250
VGGT 0.031 0.023 0.016 0.014 0.02250
Average per task 0.028 0.024 0.021 0.016
@@
CLIP ViT-B/16 0.010 0.008 0.007 0.007 0.00800
DINO ViT-B/16 0.014 0.013 0.012 0.010 0.01225
DINOv2 ViT-B/14 0.009 0.008 0.008 0.007 0.00800
DINOv3 ViT-B/16 0.005 0.006 0.004 0.003 0.0045
C-RADIOv2 ViT-B/16-CPE 0.024 0.021 0.020 0.015 0.02000
SigLIP 2 B/16-512 0.023 0.020 0.019 0.015 0.01925
TIPS ViT-B/14-HR 0.023 0.019 0.017 0.009 0.01700
VGGT 0.022 0.021 0.016 0.009 0.0170
Average per task 0.016 0.015 0.0129 0.0094
@@
CLIP ViT-B/16 0.026 0.023 0.024 0.020 0.02325
DINO ViT-B/16 0.041 0.038 0.036 0.030 0.03625
DINOv2 ViT-B/14 0.025 0.022 0.022 0.019 0.02200
DINOv3 ViT-B/16 0.016 0.015 0.009 0.005 0.01125
C-RADIOv2 ViT-B/16-CPE 0.065 0.057 0.053 0.039 0.05350
SigLIP 2 B/16-512 0.058 0.050 0.047 0.038 0.04825
TIPS ViT-B/14-HR 0.067 0.057 0.049 0.025 0.04950
VGGT 0.053 0.044 0.038 0.023 0.0395
Average per task 0.0439 0.0383 0.0348 0.0249
Table 11: System performance under memory scaling. We report wall-clock time, memory usage, CPU usage and efficiency, and memory usage and efficiency for viewpoint-based segmentation with memory sizes (a) 320k, (b) 640k, and (c) 1,024k. Some entries are marked as (running) or show zeroed usage values; these correspond to cases where monitoring logs did not update correctly, even though the jobs completed successfully. Bold values highlight the largest drops in CPU efficiency for each model.
CPU Memory
Model Wall-clock time Used Eff. Used (GB) Eff.
(a) Memory: 320k
CLIP ViT-B/16 07:32:14 3-15:33:05 16.13% 198.14 41.28%
DINO ViT-B/16 07:32:35 3-12:20:38 15.53% 214.11 44.61%
DINOv2 ViT-B/14 08:39:50 (running) 00:00:$0$ 0.00% 0.00 0.00%
DINOv3 ViT-B/16 06:48:47 00:00:03 0.00% 185.82 38.71%
C-RADIOv2 ViT-B/16-CPE 07:27:39 3-15:50:23 16.35% 199.82 41.63%
SigLIP 2 B/16-512 07:39:28 3-19:58:28 16.68% 189.89 39.56%
TIPS ViT-B/14-HR 08:33:54 (running) 00:00:$0$ 0.00% 0.00 0.00%
VGGT 10:21:59 00:00:03 0.00% 188.59 39.29%
(b) Memory: 640k
CLIP ViT-B/16 08:21:$0$ 3-10:39:36 13.75% 177.58 37.00%
DINO ViT-B/16 08:28:26 (running) 00:00:$0$ 0.00% 0.00 0.00%
DINOv2 ViT-B/14 09:43:52 (running) 00:00:$0$ 0.00% 0.00 0.00%
DINOv3 ViT-B/16 07:23:44 00:00:03 0.00% 199.28 41.52%
C-RADIOv2 ViT-B/16-CPE 08:18:48 3-11:02:33 13.87% 203.63 42.42%
SigLIP 2 B/16-512 08:26:16 3-16:01:55 14.49% 181.39 37.79%
TIPS ViT-B/14-HR 09:33:10 (running) 3-09:16:27 11.82% 209.87 43.72%
VGGT 13:00:18 00:00:03 0.00% 167.43 34.88%
(c) Memory: 1,024k
CLIP ViT-B/16 09:20:31 3-07:03:04 11.75% 207.41 43.21%
DINO ViT-B/16 09:29:42 3-07:07:59 11.58% 213.23 44.42%
DINOv2 ViT-B/14 11:10:37 3-11:36:03 10.39% 209.23 43.59%
DINOv3 ViT-B/16 08:22:03 00:00:03 0.00% 195.13 40.65%
C-RADIOv2 ViT-B/16-CPE 09:28:53 3-13:29:38 12.52% 200.34 41.74%
SigLIP 2 B/16-512 09:36:48 (running) 00:00:$0$ 0.00% 0.00 0.00%
TIPS ViT-B/14-HR 11:04:09 3-11:57:18 10.53% 211.38 44.04%
VGGT 16:07:11 00:00:03 0.00% 167.04 34.80%
E.5 Qualitative analysis
@@
Figure 29: Overlay comparison of ground truth and prediction. The same instance as in Figure 30 is shown for better. Left: ground truth overlaid on the input image. Center: prediction overlaid on the input image. Right: difference visualization, with the ground truth overlaid on the predicted mask.

### v2

@@ Appendix E Reproduction results
Appendix F Additional results
F.1 Experiment A
In experiment A, we assess model generalization across view-angle difficulties using per-class performance curves. In the figures below, the bars represent mean mIoU for unseen validation bins, and dots mark reference viewpoints. Performance typically declines as viewpoint distance increases from the reference bins, consistent with the trend in Figure 2. Across viewpoint difficulties, DINOv3, DINOv2, DINO, and CLIP consistently outperform all other models. C-RADIOv2 follows closely, while SigLIP2 and especially TIPS degrade substantially under extreme viewpoints (with TIPS showing the weakest overall robustness). DINOv3 generally outperforms all models; however, this does not hold for the coat rack and guitar stand classes, where, in the Extreme difficulty scenario, it becomes one of the worst-performing models.
Across all models, the background class performs best; other classes that perform well are durian, broccoli, and strings. These tend to generalize above average under unseen viewpoints. In contrast, bed, coat rack, guitar stand, and sink perform worst (the largest negative generalization gaps). For the sink class, DINOv3 significantly outperforms all other models. For bed, the drop is explained by poor ground-truth annotations that likely reduce model scores (as explored in Appendix F.1.1).
Figure 5: Experiment A: background. Viewpoint generalization for the background class across 4 training difficulty settings: (a) easy, (b) medium, (c) hard, and (d) extreme. All models achieve near-perfect consistency across bins, indicating that the background segmentation task remains invariant to viewpoint changes.
@@ F.1 Experiment A
Figure 7: Experiment A: bed. The bed class scores far below the all-class baseline, with it’s mIoU scores remaining low even for strong models. This degradation is largely explained by systematically incorrect ground-truth masks, as seen in Figures 2226 we analyze in Appendix F.1.1. The figure reflects this annotation-driven performance drop.
Figure 8: Experiment A: toy cat. The toy cat performance is very similar to the dataset-wide average observed in Figure 2, with relative differences near zero. Strong models (CLIP, DINO v1-v3) achieve high mIoU (0.77–0.84), and accuracy drops consistently across difficulty settings.
@@
Figure 10: Experiment A: microwave. This class underperforms relative to Figure 2, with the largest drops appearing under Hard and Extreme difficulty. Model behavior diverges notably: DINOv3, DINOv2 and CLIP maintain comparatively strong performance, whereas DINO shows a mild decline. TIPS and SigLIP2 appear to struggle when segmenting the microwave category and degrade sharply under the Extreme condition, with a pronounced collapse beyond the 45 bin. Despite these weaker generalization patterns, the overall degradation across viewpoint bins follows the expected trend.
@@
Figure 12: Experiment A: toy dragon. This class exhibits mildly negative gaps, in particular C-RADIOv2 is always lower than TIPS (when comparing to Figure 2). As a whole though, its performance trends mirror the general degradation observed of weaker models struggling as viewpoint distance increases.
Figure 13: Experiment A: coat rack. The coat rack is the worst-performing class overall, with mIoU remaining below 0.7 across all validation bins. As seen in Figure 4, the object’s structure is extremely thin, which likely makes it highly sensitive to viewpoint changes and contributes to its weak generalization. It remains an open question whether this difficulty is intrinsic to the geometry or arises from dataset-specific factors. This could be an area of further investigation.
Figure 14: Experiment A: guitar stand. The Guitar stand class scores well below the baseline for all models and difficulties. Like the coat rack, the thin structure likely contributes to poor performance. However, the exact cause remains unclear and is a point for future investigation .
@@
Figure 16: Experiment A: toilet. The toilet class consistently performs above the overall trend in Figure 2, with C-RADIOv2 showing the strongest boost, achieving increases of nearly 0.2 mIoU across difficulty levels.
Figure 17: Experiment A: sink. The sink class is one of the three worst-performing classes in the dataset. DINOv3 is the only model that performs well in this category. All other models show large negative gaps relative to Figure 2, with SigLIP2 already underperforming in the easy setting by roughly 0.2 mIoU. Absolute scores generally remain low across all difficulty levels, and the degradation with viewpoint variation is steep. What geometric or appearance factors make sinks perform poorly is not immediately obvious, suggesting that sink segmentation is a broader failure mode worth investigating.
Figure 18: Experiment A: strings. The strings class generalizes reliably across all models, with performance declining smoothly as viewpoint difficulty increases, yet remaining consistently above the dataset average. Since this category spans multiple instruments (e.g. guitar, ukulele, guzheng), future work could investigate which visual attributes they share, which lead to strong viewpoint stability.
@@
Figure 20: Experiment A: durian. Durian is among the best-performing classes, showing high absolute mIoU (0.78–0.94) across most models and all difficulty levels. This robustness is somewhat surprising: although the durian has a rich surface texture, its overall geometry is nearly spherical. It remains unclear whether the stable segmentation under viewpoint shifts stems from its distinctive visual signature or from contextual cues in the surrounding scene, such as the curtains highlighted in Figure 4.
F.1.1 Bed category ground truth analysis
@@
F.2 Experiment B
As shown in Figure 3 and Figure 27, most models exhibit a gradual performance decline as the distance of the validation bin from the training bin (0) increases. Visually, only C-RADIOv2 (ViT-B/16-CPE) and TIPS (ViT-B/14-HR) exhibit sharp performance drops with increasing validation angles, indicating limited generalization capacity under extreme viewpoint shifts.
Figure 27: Absolute mIoU under viewpoint shifts. Each curve shows the mIoU of a model evaluated on increasing validation bins, measured directly (without normalization to the 0 baseline). The horizontal axis represents viewpoint angle relative to the training bin (0), and the vertical axis shows the segmentation performance via mIoU. All models show a gradual decrease as the angular distance grows, reflecting reduced correspondence between memory and query views. DINOv2 has the most consistent generalization across bins, maintaining a high mIoU across all bins. DINOv3 maintains the highest mIoU. By contrast, TIPS displays the steepest decline (dropping by more than 0.2 between 15 and 75), confirming its reduced robustness under large viewpoint changes.
F.3 Experiment C
In Experiment C, we evaluate the impact of memory bank size on model performance. Table 9 reports the absolute mIoU gains per difficulty level when increasing the memory bank size (a) from 320k to 640k, (b) from 640k to 1,024k, and (c) from 320k to 1,024k. This highlights how the sensitivity to additional memory varies across model architectures, while also showing diminishing returns as memory size increases.
To put these performance gains into practical context, we summarize the computational setup of Experiment C. As mentioned in Appendix D, all experiments were executed on a single NVIDIA A100 GPU node with 4 GPUs. Across all configurations, runs took approximately 7–11 hours to complete, with larger memory banks consistently leading to longer runtimes. At larger memory sizes, runs encountered out-of-memory (OOM) errors unless additional measures were applied. We therefore fixed the batch size to 4 and enabled FAISS index sharding, reflecting the higher computational demands of larger memory configurations.
Table 9: Performance gains from memory. We report absolute mIoU improvements when increasing memory from (a) 320k to 640k, (b) 640k to 1,024k, and (c) 320k to 1,024k. Values are computed as differences from Table 4, with the final column showing the average gain across all 4 difficulty levels.
@@ F.3 Experiment C
CLIP ViT-B/16 0.017 0.016 0.017 0.013 0.016
DINO ViT-B/16 0.027 0.025 0.024 0.019 0.024
DINOv2 ViT-B/14 0.017 0.014 0.015 0.012 0.014
DINOv3 ViT-B/16 0.011 0.010 0.007 0.003 0.008
C-RADIOv2 ViT-B/16-CPE 0.041 0.036 0.033 0.024 0.034
SigLIP2 B/16-512 0.035 0.030 0.029 0.023 0.029
TIPS ViT-B/14-HR 0.044 0.038 0.032 0.016 0.033
Average per task 0.027 0.024 0.022 0.016 0.022
@@
CLIP ViT-B/16 0.010 0.008 0.007 0.007 0.008
DINO ViT-B/16 0.014 0.013 0.012 0.010 0.012
DINOv2 ViT-B/14 0.009 0.008 0.008 0.007 0.008
DINOv3 ViT-B/16 0.006 0.005 0.003 0.002 0.004
C-RADIOv2 ViT-B/16-CPE 0.024 0.021 0.019 0.015 0.020
SigLIP2 B/16-512 0.023 0.020 0.019 0.015 0.019
TIPS ViT-B/14-HR 0.023 0.019 0.017 0.009 0.017
Average per task 0.015 0.014 0.012 0.009 0.013
@@
CLIP ViT-B/16 0.026 0.024 0.024 0.020 0.024
DINO ViT-B/16 0.041 0.038 0.036 0.030 0.036
DINOv2 ViT-B/14 0.025 0.022 0.023 0.019 0.022
DINOv3 ViT-B/16 0.016 0.015 0.010 0.005 0.012
C-RADIOv2 ViT-B/16-CPE 0.065 0.058 0.053 0.039 0.054
SigLIP2 B/16-512 0.058 0.050 0.047 0.038 0.048
TIPS ViT-B/14-HR 0.067 0.057 0.049 0.025 0.049
Average per task 0.043 0.038 0.035 0.025 0.035
Appendix G Qualitative analysis
@@
Figure 29: Overlay comparison of ground truth and prediction. The same instance as in Figure 28 is shown for a better comparison. Left: ground truth overlaid on the input image. Center: prediction overlaid on the input image. Right: difference visualization, with the ground truth overlaid on the predicted mask.